Solve by the elimination method.
8x-y=69
x+9y=36

Answer by nerdybill(7384)

Solve by the elimination method.
8x-y=69
x+9y=36
.
Multiply the top equation by 9 and combine both equations:
72x-9y=621
x+9y= 36
----------
73x =657
x = 9
.
Plug above into:
x+9y=36
9+9y=36
9y=27
y = 3
.
solution:
x=9 and y=3